# NAMO-LLM: Efficient Navigation Among Movable Obstacles with Large Language Model Guidance

### v1

@@ Abstract
Several planners have been proposed to compute robot paths that reach desired goal regions while avoiding obstacles. However, these methods fail when all pathways to the goal are blocked. In such cases, the robot must reason about how to reconfigure the environment to access task-relevant regions—a problem known as Navigation Among Movable Objects (NAMO). While various solutions to this problem have been developed, they often struggle to scale to highly cluttered environments. To address this, we propose NAMO-LLM, a sampling-based planner that searches over robot and obstacle configurations to compute feasible plans specifying which obstacles to move, where, and in what order. Its key novelty is a non-uniform sampling strategy guided by Large Language Models (LLMs) biasing the tree construction toward directions more likely to yield a solution. We show that NAMO-LLM is probabilistically complete and demonstrate through experiments that it efficiently scales to cluttered environments, outperforming related works in both runtime and plan quality.
@@ I Introduction
Several planning algorithms have been developed to compute robot paths from an initial configuration to a desired one while avoiding obstacles [1, 2, 3, 4, 5]. However, in highly cluttered environments, obstacle-free paths to goal regions may not exist. In such cases, rather than reporting task failure, a more effective approach is to enable robots to interact with their environment by repositioning movable obstacles. This requires planning algorithms capable of determining which obstacles to move, in what order, and where to reposition them to create a feasible path to the goal; see Fig. 1. This planning problem, known as Navigation Among Movable Objects (NAMO) [6], has been proven to be NP-hard [7].
@@
Several algorithms have been proposed to address NAMO problems. Early approaches relied on heuristic solutions which lacked completeness guarantees [8]. To overcome this, search-based planning methods have been introduced. These methods either provide completeness guarantees for specific subclasses of NAMO problems or sacrifice computational efficiency to ensure completeness across a broader range of scenarios [6, 9, 10, 11]. For instance, [6] addressed NAMO problems where disconnected free-space components can be connected by moving a single obstacle at a time. This formulation allows the problem to be represented as a graph structure, enabling the use of resolution-complete planners. This work was later extended to handle cases where at most $k$ obstacles must be moved, with each obstacle requiring only a single displacement [10]. A more general approach is proposed in [11] to solve non-monotone NAMO problems, where multiple obstacles may need to be repositioned—potentially more than once—to establish connectivity between disconnected free-space components. While this algorithm is complete under mild assumptions, its computational cost scales exponentially with the number of movable objects. Extensions of these approaches to account for physical interactions between objects are presented in [12, 13], while extensions to unknown environments are considered in [14, 15, 16, 17, 18, 19]. In addition to search-based methods, sampling-based planners have been developed to improve efficiency while still maintaining (probabilistic) completeness guarantees for general NAMO problems [20, 21, 22, 23]. The key idea in these approaches is to incrementally construct trees by sampling and adding nodes, where each node represents the environment’s configuration, and edges model feasible manipulation actions. However, their efficiency declines as the number of movable objects or the number of required obstacle relocations grows, due to their uniform sampling strategy.
Figure 1: A ground robot is tasked with reaching a destination behind the building block; see (1). All direct pathways are initially blocked (see Fig. 2). The robot autonomously removes obstacles to clear a path; see (2)–(4).
@@
This paper addresses this limitation by designing a new sampling-based NAMO algorithm that efficiently scales to environments characterized by large numbers of movable obstacles and/or the need for multiple obstacle relocations to reach desired regions. Inspired by [22], our algorithm incrementally builds trees that explore the space of obstacle and robot configurations. Each tree node represents an environment configuration, while edges among nodes correspond to manipulation actions that relocate movable obstacles. At each iteration, a tree node and a manipulation action are sampled. Application of that action to the sampled tree node results in new tree node that is added to the tree if feasible. The key novelty of our work lies in a non-uniform sampling strategy that leverages Large Language Models (LLMs) to accelerate plan generation. Specifically, our sampling strategy directs the tree to grow towards LLM-recommended directions at certain times while allowing random exploration otherwise. We refer to the resulting planner as NAMO-LLM, for Navigation Among Movable Obstacles with Large Language Model guidance. We show that NAMO-LLM is probabilistically complete meaning that the probability of computing a feasible plan converges to one as the number of iterations goes to infinity. We conduct extensive comparative experiments against both search-based planners [18] and sampling-based planners relying on uniform sampling strategies [22], demonstrating that our method significantly reduces both the time required to compute the first feasible plan and the corresponding number of obstacles that need to be relocated. We also demonstrate NAMO-LLM on hardware experiments; see Fig. 1.
@@
Related Works: In addition to the NAMO algorithms discussed earlier, our work is also related to research on rearrangement planning (RP) problems where robots manipulate movable obstacles [24, 25, 26, 27]. RP differs from NAMO because it considers predefined goal positions for the movable obstacles and typically does not specify a final goal for the robot. While our work primarily focuses on NAMO problems, we believe that our non-uniform sampling strategy can also be integrated with various sampling-based approaches for planning among movable obstacles, including RP and manipulation planning [28, 29]. However, a formal integration with these approaches is beyond the scope of this paper. Furthermore, several works design non-uniform sampling strategies to accelerate sampling-based planners [30, 31, 32, 33, 34, 35, 36]. For example, [34] learns a non-uniform sampling distribution from demonstrations. However, we emphasize that these works target different planning problems from ours, focusing on designing obstacle-free paths between an initial and a desired robot configuration (while also assuming that such paths exist). As such, they cannot be directly applied to our setting. It is worth noting that, unlike [32, 33, 34, 35, 36], our work does not require training neural networks; instead, ours leverages pre-trained LLMs.
@@
Summary of Contributions: First, we introduce NAMO-LLM, a novel sampling-based planner designed to efficiently solve NAMO problems in highly cluttered environments. Second, to the best of our knowledge, this is the first integration of LLMs within a sampling-based NAMO planner, significantly accelerating plan generation. Third, we conduct extensive experiments demonstrating that our approach outperforms existing planning algorithms in terms of efficiency and solution quality. Fourth, we release an open-source implementation of our proposed algorithm [37].
@@ II Problem Formulation
Robot Modeling: We consider a robot with the following kinematics: ${\mathbf{p}}(t+1)={\mathbf{f}}({\mathbf{p}}(t),{\mathbf{u}}(t)),$ where $\mathbf{p}(t)\in{\mathcal{P}}$ and $\mathbf{u}(t)\in{\mathcal{U}}$ denote the robot configuration (e.g., position and orientation) and the control input at time step $t$, respectively. We assume that ${\mathbf{p}}(t)$ is known at all times and that the robot is holonomic and, therefore, it can move along any desired path. The robot is also equipped with a gripper, enabling it to manipulate one movable object at a time.
Figure 2: Top-view representation of the urban environment shown in Fig. 1. The set ${\mathcal{P}}$ is defined by the position of the center of the disk-shaped robot. Given ${\mathcal{N}}(0)$, the set ${\mathcal{O}}_{M}(0)$ includes only obstacles 1-6.
@@
Environment Modeling: The robot operates in an environment $\Omega$ containing a set of known obstacles ${\mathcal{O}}=\{o_{1},\dots,o_{O}\}$. The obstacles are classified as movable obstacles, collected in a set ${\mathcal{O}}_{\text{movable}}\subseteq{\mathcal{O}}$, and non-movable ones. A movable object can be repositioned by the robot as long as it first grasps it. The configuration (e.g., position and orientation) of an obstacle $o_{i}\in{\mathcal{O}}$ at time $t$ is denoted by $c_{i}(t)\in{\mathcal{C}}_{i}$ where ${\mathcal{C}}_{i}$ denotes its configuration space. We also denote by $co_{i}(t)$ the region in the configuration space $\mathcal{P}$ of the robot that corresponds to configurations where the robot is in collision with obstacle $o_{i}$; see Fig. 2. Thus, the robot’s free space at time $t$, denoted by ${\mathcal{P}}_{\text{free}}(t)$, is defined as ${\mathcal{P}}_{\text{free}}(t)={\mathcal{P}}\setminus\mathcal{CO}(t)$ where $\mathcal{CO}(t)=\cup_{i=1}^{n_{O}}\{co_{i}(t)\}$. Notice that ${\mathcal{P}}_{\text{free}}(t)$ may consist of multiple connected components whose boundaries are determined by the configuration of the obstacles. We denote by ${\mathcal{N}}(t)\subseteq{\mathcal{P}}_{\text{free}}(t)$, the free-space component that ${\mathbf{p}}(t)$ lies in; see Fig. 2.
@@
Robot Task: The robot is tasked with reaching a desired configuration starting from an initial one ${\mathbf{p}}(0)$ while avoiding collisions with obstacles. Hereafter, we denote by ${\mathcal{P}}_{g}\subseteq{\mathcal{P}}$ a set collecting all desired robot configurations. We consider cases where all configurations in ${\mathcal{P}}_{g}$ are initially unreachable due to the initial configuration of obstacles, i.e., ${\mathcal{P}}_{g}\cap{\mathcal{N}}(0)=\emptyset$. In such cases, the robot must determine which movable obstacles need to be relocated and where they should be placed. To formalize this, we define the set of manipulable obstacles at time $t$ as $\mathcal{O}_{M}(t)\subseteq{\mathcal{O}}_{\text{movable}}$ which includes all movable obstacles $o_{i}$ that the robot can directly reach, i.e., $\mathcal{O}_{M}(t)=\{o_{i}\in{\mathcal{O}}_{\text{movable}}\mid co_{i}(t)\cap{\mathcal{N}}(t)\neq\emptyset\}$; see also Fig. 2. Next, we define manipulation actions of the form $a(c_{i}(t),c_{i}^{\prime})$ which involve moving a manipulable object $o_{i}\in\mathcal{O}_{M}(t)$, from its current configuration $c_{i}(t)$ to a new configuration $c_{i}(t^{\prime})=c_{i}^{\prime}$, for some $t^{\prime}>t$. In other words, the application of this action requires the robot to approach the obstacle’s current location (determined by $c_{i}(t)$), grasp it, and relocate it to the new configuration $c_{i}^{\prime}$. Thus, our goal is to compute a plan $\tau$, i.e., a sequence of manipulation actions so that eventually ${\mathcal{P}}_{g}\cap{\mathcal{N}}(t)\neq\emptyset$. We defined this plan as $\tau=\tau(0),\tau(1),\dots,\tau(k),\dots,\tau(H)$, where $\tau(k)$ refers to the $k$-th manipulation action, $k\in\{0,1,\dots,H\}$, for some horizon $H\geq 0$. Note that the index $k$ is different from the time step $t$, as the application of an action $\tau(k)$ may require multiple time steps $t$. The problem that this paper addresses can be summarized as follows:
@@
Given an initial robot configuration ${\mathbf{p}}(0)$, a set ${\mathcal{P}}_{g}$ of desired configurations, and initial obstacle configurations $c_{i}(0)$, compute a manipulation plan $\tau=\tau(0),\tau(1),\dots,\tau(k),\dots,\tau(H)$, so that ${\mathcal{P}}_{g}\cap{\mathcal{N}}(H)\neq\emptyset$.
@@ III Proposed Planning Algorithm
In this section, we present NAMO-LLM, our proposed planner to address Problem 1. In Section III-A, we present a sampling-based planner to design plans $\tau$. In Section III-B, we present a non-uniform sampling strategy that accelerates the plan synthesis phase by leveraging Large Language Models (LLM). In Section III-C, we discuss the input prompt to the LLM guiding the sampling strategy. In Section III-D, we show that NAMO-LLM is probabilistically complete.
@@ 5       Sample a node ${\mathbf{q}}_{\text{rand}}\in{\mathcal{V}}$ using $f_{\mathcal{V}}$;
6       $[{\mathbf{q}}_{\text{new}},a(c_{i}(t),c_{i}^{\prime})]=\texttt{TreeExpansion}({\mathbf{q}}_{\text{rand}})$;
@@
9             $A(({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}}))=a(c_{i}(t),c_{i}^{\prime})$;
@@ 19       Using $A$ and $\tau_{{\mathbf{q}}}$, compute the manipulation plan $\tau$ ;
20
21else
@@ 22       No plan $\tau$ has been found;
1 Compute the set of manipulable obstacles ${\mathcal{O}}_{M,\text{rand}}(t)$ ;

### v2

@@ Abstract
Several planners have been proposed to compute robot paths that reach desired goal regions while avoiding obstacles. However, these methods fail when all pathways to the goal are blocked. In such cases, the robot must reason about how to reconfigure the environment to access task-relevant regions—a problem known as Navigation Among Movable Objects (NAMO). While various solutions to this problem have been developed, they often struggle to scale to highly cluttered environments. To address this, we propose NAMO-LLM, a sampling-based planner that searches over robot and obstacle configurations to compute feasible plans specifying which obstacles to move, where, and in what order. Its key novelty is a non-uniform sampling strategy guided by Large Language Models (LLMs) biasing the search process toward directions more likely to yield a solution. We show that NAMO-LLM is probabilistically complete and demonstrate through experiments that it efficiently scales to cluttered environments, outperforming related works in both runtime and plan quality.
@@ I Introduction
Several planning algorithms have been developed to compute robot paths from an initial configuration to a desired one while avoiding obstacles [1, 2]. However, in highly cluttered environments, obstacle-free paths to goal regions may not exist; see Fig. 1. In such cases, rather than reporting task failure, a more effective approach is to enable robots to interact with their environment by relocating movable obstacles. This requires planners capable of determining which obstacles to move, in what order, and where to reposition them to create a feasible path to the goal; see Fig. 1. This problem, known as Navigation Among Movable Objects (NAMO) has been proven to be NP-hard [3].
@@
Early approaches to address NAMO problems relied on heuristic solutions which lacked completeness guarantees [4]. To overcome this, search-based planning methods have been introduced. These methods either provide completeness guarantees for specific subclasses of NAMO problems or sacrifice computational efficiency to ensure completeness across a broader range of scenarios [5, 6, 7, 8]. Extensions of these approaches to unknown environments are considered in [9, 10, 11, 12, 13]. Sampling-based planners have been developed to improve efficiency while still maintaining (probabilistic) completeness guarantees for general NAMO problems [14, 15, 16, 17]. Their key idea is to incrementally construct trees by sampling nodes, where each node represents the environment’s configuration, and edges model feasible manipulation actions. However, their efficiency declines as the number of movable obstacles or the number of required obstacle relocations grows, due to their uniform sampling strategy.
Figure 1: A robot must reach a destination behind the building block. Since all direct paths are blocked, it relocates obstacles to access the goal region.
@@
This paper addresses this limitation by designing a new sampling-based NAMO algorithm that efficiently scales to environments with many movable obstacles and/or requiring multiple obstacle relocations to reach target regions. Inspired by [16], our algorithm incrementally builds trees exploring the configuration space of movable objects and the robot’s free-space components. The key novelty lies in a non-uniform sampling strategy that leverages Large Language Models (LLMs) to accelerate plan generation. Specifically, the strategy directs the tree toward LLM-recommended directions at certain times while allowing random exploration otherwise. We refer to the resulting planner as NAMO-LLM and show it is probabilistically complete, meaning the probability of computing a feasible plan (if one exists) converges to one as the number of iterations grows. Extensive numerical and hardware experiments demonstrate that NAMO-LLM outperforms search-based planners [12], uniform-sampling planners [16], and planners relying solely on LLMs in terms of runtime for the first feasible plan, the number of obstacles relocated, and task success rates. This performance gap increases with environment complexity.
@@
Related Works: In addition to the NAMO algorithms discussed earlier, our work is also related to rearrangement planning (RP) problems where robots manipulate movable obstacles [18, 19, 20, 21]. RP differs from NAMO because it considers predefined goal positions for the movable obstacles and typically does not specify a final goal for the robot. We believe that our non-uniform sampling strategy can also be integrated with various sampling-based approaches for planning among movable obstacles, including RP and manipulation planning [22, 23]; however, this integration is beyond the scope of this paper. Furthermore, several works design non-uniform sampling strategies to accelerate sampling-based planners [24, 25, 26, 27, 28]. However, we emphasize that these works target different planning problems from ours, focusing on designing obstacle-free paths between an initial and a desired robot configuration (while also assuming that such paths exist). As such, they cannot be directly applied to our setting. It is worth noting that, unlike [25, 26, 27, 28], our work does not require training neural networks; instead, ours leverages pre-trained LLMs.
@@
Summary of Contributions: First, we introduce NAMO-LLM, a novel sampling-based planner designed to efficiently solve NAMO problems in highly cluttered environments. Second, we propose the first integration of LLMs within a sampling-based NAMO planner, significantly accelerating plan generation. Third, we conduct extensive experiments demonstrating that our approach scales well to complex environments and outperforms related planners in efficiency, solution quality, and task success rates.
@@ II Problem Formulation
Robot Modeling: We consider a robot modeled as: ${\mathbf{p}}(t+1)={\mathbf{f}}({\mathbf{p}}(t),{\mathbf{u}}(t)),$ where $\mathbf{p}(t)\in{\mathcal{P}}$ and $\mathbf{u}(t)\in{\mathcal{U}}$ denote the robot configuration (e.g., position and orientation) and the control input at time step $t$, respectively. We assume that ${\mathbf{p}}(t)$ is always known and that the robot is holonomic and, therefore, it can move along any desired path. The robot is also equipped with a gripper, enabling it to manipulate one movable object at a time.
@@
Environment Modeling: The robot resides in an environment $\Omega$ containing a set of known obstacles ${\mathcal{O}}=\{o_{1},\dots,o_{O}\}$. The obstacles are classified as movable obstacles, collected in a set ${\mathcal{O}}_{\text{movable}}\subseteq{\mathcal{O}}$, and non-movable ones. A movable object can be repositioned by the robot as long as it first grasps it. The configuration (e.g., position and orientation) of an obstacle $o_{i}\in{\mathcal{O}}$ at time $t$ is denoted by $c_{i}(t)\in{\mathcal{C}}_{i}$ where ${\mathcal{C}}_{i}$ denotes its configuration space. We also denote by ${\bf{co}}_{i}(t)$ the region in $\mathcal{P}$ that corresponds to configurations where the robot is in collision with obstacle $o_{i}$. Thus, the robot’s free space at time $t$, denoted by ${\mathcal{P}}_{\text{free}}(t)$, is defined as ${\mathcal{P}}_{\text{free}}(t)={\mathcal{P}}\setminus\mathcal{CO}(t)$ where $\mathcal{CO}(t)=\cup_{i=1}^{n}\{{\bf{co}}_{i}(t)\}$. Notice that ${\mathcal{P}}_{\text{free}}(t)$ may consist of multiple connected components whose boundaries are determined by the configuration of the obstacles. We denote by ${\mathcal{N}}(t)\subseteq{\mathcal{P}}_{\text{free}}(t)$, the free-space component that ${\mathbf{p}}(t)$ lies in; see also Fig. 2.
@@
Robot Task: The robot must reach a desired configuration from an initial one ${\mathbf{p}}(0)$ while avoiding obstacles. Let ${\mathcal{P}}_{g}\subseteq{\mathcal{P}}$ denote the desired configurations. We focus on cases where all configurations in ${\mathcal{P}}_{g}$ are initially unreachable due to obstacle placements, i.e., ${\mathcal{P}}_{g}\cap{\mathcal{N}}(0)=\emptyset$. The robot must then determine which movable obstacles to relocate and their target placements. To formalize this, we define the set of manipulable obstacles at time $t$ as $\mathcal{O}_{M}(t)\subseteq{\mathcal{O}}_{\text{movable}}$ which includes all movable obstacles $o_{i}$ that the robot can directly reach, i.e., $\mathcal{O}_{M}(t)=\{o_{i}\in{\mathcal{O}}_{\text{movable}}\mid{\bf{co}}_{i}(t)\cap{\mathcal{N}}(t)\neq\emptyset\}$. Next, we define manipulation actions of the form $a(c_{i}(t),c_{i}^{\prime})$ requiring the robot to approach, grasp, and relocate an an object $o_{i}\in\mathcal{O}_{M}(t)$, from its current configuration $c_{i}(t)$ to a new configuration $c_{i}(t^{\prime})=c_{i}^{\prime}$, for some $t^{\prime}>t$. Our goal is to compute a plan $\tau=\tau(0),\dots,\tau(H)$ such that ${\mathcal{P}}_{g}\cap{\mathcal{N}}(H)\neq\emptyset$ for some horizon $H\geq 0$. Here, $\tau(k)$ denotes the $k$-th manipulation action, with $k\in\{0,1,\dots,H\}$. Note that $k$ indexes actions, not time steps, as executing $\tau(k)$ may span multiple time steps $t$.
@@
Given ${\mathbf{p}}(0)$, ${\mathcal{P}}_{g}$, $\mathcal{O}$}, ${\mathcal{O}}_{\text{movable}}$, and $c_{i}(0)$ for all $o_{i}$, compute $\tau=\tau(0),\dots,\tau(H)$ so that ${\mathcal{P}}_{g}\cap{\mathcal{N}}(H)\neq\emptyset$.
@@ III Proposed Planning Algorithm
In this section, we present NAMO-LLM, our planner for Problem 1. Section III-A describes a sampling-based planner for generating plans $\tau$. Section III-B introduces a non-uniform sampling strategy leveraging LLMs to accelerate plan synthesis. In Section III-C, we discuss the input prompt to the LLM guiding the sampling strategy. In Section III-D, we show that NAMO-LLM is probabilistically complete.
Figure 2: Tree structure built by the proposed planner. Each node (blue) represents a state ${\mathbf{q}}(t)$, and each edge (green arrow) represents a manipulation action. For instance, ${\mathbf{q}}_{\text{new}}$ is reached from ${\mathbf{q}}_{\text{rand}}$ via action $a({\mathbf{c}}_{5},{\mathbf{c}}_{5}^{\prime})$ moving obstacle $o_{5}$. If moving $o_{5}$ is recommended by the LLM (rather than random sampling), the score $v({\mathbf{q}}_{\text{new}})$ is set to 11; otherwise, it remains 0. The label next to each node indicates whether it was generated by moving an LLM-suggested obstacle (L) or a randomly selected obstacle (R) (see Section III-B). The number next to the label is the score $u({\mathbf{q}})$ yielding a set ${\mathcal{V}}_{v}$ that collects the top two gray-shaded nodes (see Section III-B). The two snapshots illustrate the environment associated with ${\mathbf{q}}_{\text{rand}}$ and ${\mathbf{q}}_{\text{new}}$.
III-A Sampling-based Planning
We propose a sampling-based planner to address Problem 1, summarized in Alg. 1. The algorithm incrementally builds a tree exploring the configuration space of the movable objects and the robot’s free-space components to compute $\tau$; see Fig. 2. We denote the constructed tree by ${\mathcal{G}}=\{{\mathcal{V}},{\mathcal{E}}\}$, where $\mathcal{V}$} and $\mathcal{E}$} is a set of nodes and edges, respectively. The set $\mathcal{V}$} contains states of the form ${\mathbf{q}}(t)=[\mathcal{N}(t),{\mathbf{c}}(t)]$, where ${\mathbf{c}}(t)$ is a vector stacking the configuration of all obstacles, i.e., ${\mathbf{c}}(t)=[{\mathbf{c}}_{1}(t),{\mathbf{c}}_{2}(t),\dots,{\mathbf{c}}_{n}(t)]$ and ${\mathcal{N}}(t)$ denotes the free-space component that ${\mathbf{p}}(t)$ lies in. The root of the tree models the initial configuration of the obstacles and the robot, i.e., ${\mathbf{q}}(0)=[{\mathcal{N}}(0),{\mathbf{c}}(0)]$ [lines 1-1, Alg. 1]. The tree is initialized so that ${\mathcal{V}}=\{{\mathbf{q}}(0)\}$ and $\mathcal{E}$$\emptyset$ [line 1, Alg. 1]. Also, the tree is built incrementally by sampling new states ${\mathbf{q}}_{\text{new}}$ that are added to the tree structure [line 1, Alg. 1]. The algorithm terminates after a user-specified number of iterations $m_{\text{max}}$ and returns a feasible plan $\tau$ (if it has been found). In what follows, we present our method in detail.
@@ 5  Sample a node ${\mathbf{q}}_{\text{rand}}\in{\mathcal{V}}$ using $f_{\mathcal{V}}$;
6  ${\mathbf{q}}_{\text{new}},a({\mathbf{c}}_{i}(t),{\mathbf{c}}_{i}^{\prime})]=\texttt{TreeExpansion}({\mathbf{q}}_{\text{rand}})$;
@@
9     $A(({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}}))=a({\mathbf{c}}_{i}(t),{\mathbf{c}}_{i}^{\prime})$;
@@ 19  Using $A$ and $\tau_{{\mathbf{q}}}$, compute the manipulation plan $\tau$ ;
20  Compute robot motion path $\xi$ executing $\tau$;
21
22else
@@ 23  No plan $\tau$ has been found;
1Compute the set of manipulable obstacles ${\mathcal{O}}_{M,\text{rand}}(t)$ ;